An FM station broadcasts classical music at 94.7 MHz (megahertz, or 106 Hz). Find the wavelength (in m, nm, and Å) of these radio waves.

_______ m

nm (Enter your answer in scientific notation.)

____Å (Enter your answer in scientific notation.)
Answer #1

f = 94.7 MHz
= 94.7*10^6 Hz
= 9.47*10^7 Hz


use:
lambda = c/f
=(3.0*10^8 m/s)/(9.47*10^7 s-1)
= 3.17 m

Answer:
3.17 m
3.17*10^9 nm
3.17*10^10 A